Solve for x (Hint: Factor the left-hand side)

x^2 + 4x + 4 =7

Thank You!!!

Answer by stanbon(75887)

x^2 + 4x + 4 =7
(x+2)^2 = 7
Take the square root of both sides to get:
x+2=sqrt7 or x+2=-sqrt7
x=-2+sqrt7 or x=-2-sqrt7
Cheers,
Stan H.